# Solving Fredholm Integral Equations of the Second Kind via Wasserstein Gradient Flows

### v1

@@ C.2 Proof of Proposition 3
Then, we use the following version of Young’s inequality, $\alpha\beta\leq\alpha^{2}/(2\epsilon)+\beta^{2}\epsilon/2$ for all $\epsilon>0$, to show that
@@
As shown in (33), the first component of the drift of the MKVSDE (9) is Lipschitz continuous. In addition, using the fact that under Assumption 1 $\mathrm{k},\varphi\in\mathrm{C}^{\infty}(\mathbb{R}^{d}\times\mathbb{R}^{d},[0,+\infty))$ and Leibniz integral rule for differentiation under the integral sign (e.g. [8, Theorem 16.8]), we have that $b(x,\nu)$ in (29) is $\mathrm{C}^{\infty}(\mathbb{R}^{d},\mathbb{R}^{d})$ for all fixed $\nu\in\mathcal{P}(\mathbb{R}^{d})$. Finally, we need to show that
$\nabla b(x,\nu)$ $:=\int\limits_{\mathbb{R}^{d}}\nabla_{1}b^{\eta}(x,z,\nu)\mathrm{d}\nu(z)=\nabla b^{1}(x,\nu)+\nabla b^{2}(x,\nu),$
@@
where we used the fact that $\mathrm{k}$ is Lipschitz continuous with Lipschitz constant $\mathtt{M}$ and, for the last term, the dual representation of the $\mathbf{W}_{1}$ distance. In addition, since $\mathrm{k}\in\mathrm{C}^{\infty}(\mathbb{R}^{d}\times\mathbb{R}^{d},[0,+\infty))$ and $\left\|\mathrm{d}^{3}\mathrm{k}(x,y)\right\|\leq\mathtt{M}$, the function
$z\mapsto\frac{\lambda\nabla_{2}^{2}\mathrm{k}(z,x_{n})}{\lambda\nu_{n}\left[\mathrm{k}(z,\cdot)\right]+\varphi(z)+\eta}$
@@
for some finite $C$.
Since $\lim_{n\to+\infty}(x_{n},\nu_{n})=(x,\nu)$ implies $\left\|x-x_{n}\right\|\to 0$ and $\mathbf{W}_{1}(\nu_{n},\nu)\to 0$, we have that $\nabla b(x,\nu)$ is jointly continuous. Then, the result follows directly from [30, Theorem 2.11].
Appendix D Particle System and Time Discretization
D.1 Proof of Proposition 4
For any $\{x_{1}^{k,N}\}_{k=1}^{N},\{x_{2}^{k,N}\}_{k=1}^{N}\in(\mathbb{R}^{d})^{N}$ and $\ell\in\{1,\dots,N\}$, let us denote $\pi_{1}^{N}:=(1/N)\sum_{k=1}^{N}\delta_{x_{1}^{k,N}}$ and $\pi_{2}^{N}:=(1/N)\sum_{k=1}^{N}\delta_{x_{2}^{k,N}}$. Using (34) and the fact that for empirical measures
$\mathbf{W}_{1}(\pi_{1}^{N},\pi_{2}^{N})\leq N^{-1/2}\left\|x_{1}^{1:N}-x_{2}^{1:N}\right\|$
we have that
@@ D.1 Proof of Proposition 4
The existence and strong uniqueness of a solution to (11) is then a straightforward consequence of the above and [36, Chapter 5, Theorem 2.9 and 2.5].
@@
$\qquad\leq\mathtt{C}_{1}(\left\|x_{1}-x_{2}\right\|+|\pi_{2}\left[\mathrm{k}(x_{2},\cdot)\right]-\pi_{1}\left[\mathrm{k}(x_{1},\cdot)\right]|)+\mathtt{C}_{2}\mathbf{W}_{1}(\pi_{1},\pi_{2})+\alpha\mathtt{L}\left\|x_{1}-x_{2}\right\|$
$\qquad\leq(\mathtt{C}_{1}+\mathtt{C}_{2}+\alpha\mathtt{L})(\left\|x_{1}-x_{2}\right\|+|\pi_{2}\left[\mathrm{k}(x_{2},\cdot)\right]-\pi_{1}\left[\mathrm{k}(x_{1},\cdot)\right]|+\mathbf{W}_{1}(\pi_{1},\pi_{2})|).$
@@
$\qquad\leq\int\limits_{0}^{t}\mathbb{E}\left[\left\|b(\bm{X}_{s},\pi_{s})-b(\bm{X}_{s}^{1,N},\pi^{N}_{s})\right\|\right]\mathrm{d}s$
@@
$\qquad+(\mathtt{C}_{1}+\mathtt{C}_{2}+\alpha\mathtt{L})\int\limits_{0}^{t}\mathbb{E}\left[|\frac{1}{N}\sum_{i=1}^{N}\mathrm{k}(\bm{X}_{s}^{1,N},\bm{X}_{s}^{i,N})-\pi_{s}\left[\mathrm{k}(\bm{X}_{s},\cdot)\right]|\right]\mathrm{d}s$
$\qquad+(\mathtt{C}_{1}+\mathtt{C}_{2}+\alpha\mathtt{L})\int\limits_{0}^{t}\mathbf{W}_{1}(\pi_{s},\pi_{s}^{N})\mathrm{d}s$
@@
$\qquad+(\mathtt{C}_{1}+\mathtt{C}_{2}+\alpha\mathtt{L})\int\limits_{0}^{t}\mathbb{E}\left[|\frac{1}{N}\sum_{i=1}^{N}\mathrm{k}(\bm{X}_{s}^{1,N},\bm{X}_{s}^{i,N})-\pi_{s}\left[\mathrm{k}(\bm{X}_{s},\cdot)\right]|\right]\mathrm{d}s,$
where the last inequality follows from the fact that
$\mathbf{W}_{1}(\pi_{s},\pi_{s}^{N})\leq\mathbf{W}_{2}(\pi_{s},\pi_{s}^{N})\leq\left\|\bm{X}_{s}-\bm{X}_{s}^{1,N}\right\|.$ (37)
@@
where we used the Lipschitz continuity of $\mathrm{k}$, (37) and the fact that $\{(\bm{X}_{t}^{k,N})_{t\geq 0}\}_{k=1}^{N}$ is exchangeable to obtain the last inequality. Plugging the above into (36) and using Jensen’s inequality we obtain
$+\frac{\mathtt{C}_{1}+\mathtt{C}_{2}+\alpha\mathtt{L}}{N}\int_{0}^{t}\mathbb{E}\left[\sum_{i=1}^{N}|\mathrm{k}(\bm{X}_{s}^{1,N},\bm{X}_{s}^{i,\star})-\pi_{s}\left[\mathrm{k}(\bm{X}_{s},\cdot)\right]|^{2}\right]^{1/2}\mathrm{d}s.$
@@
$\mathbb{E}\left[|\mathrm{k}(\bm{X}_{s}^{1,N},\bm{X}_{s}^{i,\star})-\pi_{s}\left[\mathrm{k}(\bm{X}_{s},\cdot)\right]|^{2}\mid\bm{X}_{s}^{1,N}\right]$ $=\textrm{var}\left(\mathrm{k}(\bm{X}_{s}^{1,N},\bm{X}_{s}^{i,\star})\mid\bm{X}_{s}^{1,N}\right)$
@@ D.2 Proof of Proposition 5
with $b^{\eta}$ defined in (10). Recall that under Assumption 1 and 2 we have (34), i.e., there exist $C_{1}\geq 0$ such that for any $x_{1},x_{2}\in\mathbb{R}^{d}$ and $\pi_{1},\pi_{2}\in\mathcal{P}(\mathbb{R}^{d})$
@@
$\leq\lambda\mathtt{M}/\eta+\lambda(\mathtt{M}+1)/\eta:=C_{0}.$
@@
$\qquad\leq 2C_{1}N^{1/2}\left\|x_{1}^{1:N}-x_{2}^{1:N}\right\|.$ (42)
@@
$\qquad\leq-\alpha\mathtt{m}\left\|x_{1}^{1:N}-x_{2}^{1:N}\right\|^{2}+\alpha\mathtt{c}N+2C_{0}N^{1/2}\left\|x_{1}^{1:N}-x_{2}^{1:N}\right\|.$ (44)
Let $R=\max(4C_{0}N^{1/2}/(\alpha\mathtt{m}),(2\mathtt{c}N/\mathtt{m})^{1/2})$ with $\mathtt{m}$ and $\mathtt{c}$ as in Assumption 2(c). Then, for any $x^{1:N}_{1},x_{2}^{1:N}\in(\mathbb{R}^{d})^{N}$ with $\left\|x_{1}^{1:N}-x_{2}^{1:N}\right\|\geq R$ we further have
$\langle B_{N}(x^{1:N}_{1})-B_{N}(x^{1:N}_{2}),x_{1}^{1:N}-x_{2}^{1:N}\rangle$ $\leq-(\alpha\mathtt{m}/2)\left\|x_{1}^{1:N}-x_{2}^{1:N}\right\|^{2}.$ (45)
We conclude upon combining (41), (43) and [19, Corollary 2].
D.3 Proof of Proposition 6
@@
$\bm{M}_{t}^{N}:=(1/2)\left\|\bm{X}_{t}^{1,N}\right\|^{2}-\int\limits_{0}^{t}\left\{\langle\bm{X}_{u}^{1,N},b(\bm{X}_{u}^{1,N},\pi^{N}_{u})\rangle+\alpha d\right\}\mathrm{d}u,$
where $\{(\bm{X}_{t}^{k,N})_{t\geq 0}\}_{k=1}^{N}$ is given in (11). Using Itô’s formula we have that
$(1/2)\left\|\bm{X}_{t}^{1,N}\right\|^{2}=(1/2)\left\|\bm{X}_{s}^{1,N}\right\|^{2}+\int\limits_{s}^{t}\langle\bm{X}_{u}^{1,N},b(\bm{X}_{u}^{1,N},\pi^{N}_{u})\rangle\mathrm{d}u+d\alpha(t-s).$
Denote by $(\mathcal{F}_{t}^{N})_{t\geq 0}$ the natural filtration of $\{(\bm{B}_{t}^{k})_{t\geq 0}\}_{k=1}^{N}$. Then, $\mathbb{E}\left[\bm{M}_{t}^{N}\mid\mathcal{F}_{s}^{N}\right]=\bm{M}_{s}^{N}$, showing that $\bm{M}_{t}^{N}$ is an $\mathcal{F}^{N}$-local martingale.
@@
$(1/2)\mathbb{E}\left[\left\|\bm{X}_{t}^{1,N}\right\|^{2}\right]-(1/2)\mathbb{E}\left[\left\|\bm{X}_{s}^{1,N}\right\|^{2}\right]$
$\qquad=\mathbb{E}\left[\int\limits_{s}^{t}\left\{\langle\bm{X}_{u}^{1,N},b(\bm{X}_{u}^{1,N},\pi^{N}_{u})\rangle+\alpha d\right\}\mathrm{d}u\right]$
$\qquad=\mathbb{E}\left[\int\limits_{s}^{t}\left\{\langle\bm{X}_{u}^{1,N},\int\limits_{\mathbb{R}^{d}}b^{\eta}(\bm{X}_{u}^{1,N},z,\pi^{N}_{u})\mathrm{d}\pi_{u}^{N}(z)\rangle\right\}\mathrm{d}u\right]$
$\qquad-\mathbb{E}\left[\int\limits_{s}^{t}\left\{\langle\bm{X}_{u}^{1,N},\alpha\nabla U(\bm{X}_{u}^{1,N})\rangle-\alpha d\right\}\mathrm{d}u\right],$
@@
$(1/2)\mathbb{E}\left[\left\|\bm{X}_{t}^{1,N}\right\|^{2}\right]-(1/2)\mathbb{E}\left[\left\|\bm{X}_{s}^{1,N}\right\|^{2}\right]$
@@
$\qquad+\mathbb{E}\left[\int\limits_{s}^{t}\left\{-\alpha\mathtt{m}\left\|\bm{X}_{u}^{1,N}\right\|^{2}+\alpha\left\|\nabla U(0)\right\|\left\|\bm{X}_{u}^{1,N}\right\|+\alpha\mathtt{c}+\alpha d\right\}\mathrm{d}u\right]$
@@
$\qquad+\int\limits_{s}^{t}\left\{-\alpha\mathtt{m}\mathbb{E}\left[\left\|\bm{X}_{u}^{1,N}\right\|^{2}\right]+\alpha\left\|\nabla U(0)\right\|\mathbb{E}\left[\left\|\bm{X}_{u}^{1,N}\right\|\right]+\alpha\mathtt{c}+\alpha d\right\}\mathrm{d}u,$
@@
Let $\mathcal{V}_{t}^{N}=(1/2)\mathbb{E}\left[\left\|\bm{X}_{t}^{1,N}\right\|^{2}\right]$. Appealing to the fundamental theorem of calculus we get that
$(1/2)\mathrm{d}\mathcal{V}_{t}^{N}/\mathrm{d}t\leq$ $\mathbb{E}\left[\int\limits_{\mathbb{R}^{d}}\left\|b^{\eta}(\bm{X}_{t}^{1,N},z,\pi^{N}_{t})\right\|\left\|\bm{X}_{t}^{1,N}\right\|\mathrm{d}\pi_{t}^{N}(z)\right]$
$-\alpha\mathtt{m}\mathbb{E}\left[\left\|\bm{X}_{t}^{1,N}\right\|^{2}\right]+\alpha\left\|\nabla U(0)\right\|\mathbb{E}\left[\left\|\bm{X}_{t}^{1,N}\right\|\right]+\alpha\mathtt{c}+\alpha d$

### v2

@@ D.2 Proof of Proposition 4
Then, using Young’s inequality, $\alpha\beta\leq\alpha^{2}/(2\epsilon)+\beta^{2}\epsilon/2$ for all $\epsilon>0$,
@@
The intrinsic derivative of $\mathcal{F}^{\eta}$ is given by the first component of the drift $b$ in (13):
$D_{\nu}\mathcal{F}^{\eta}(x,\nu):=\int b^{\eta}(x,z,\nu)\mathrm{d}\nu(z)$
with $b^{\eta}$ in (10). As shown in (43), $D_{\nu}\mathcal{F}^{\eta}$ is Lipschitz continuous and thus $\mathcal{F}^{\eta}$ is continuously differentiable. In addition, using the fact that under Assumption 1 $\mathrm{k}\in\mathrm{C}^{\infty}(\mathbb{R}^{d}\times\mathbb{R}^{d},[0,+\infty))$ and Leibniz integral rule for differentiation under the integral sign (e.g. [8, Theorem 16.8]), we have that $D_{\nu}\mathcal{F}^{\eta}$ is $\mathrm{C}^{\infty}(\mathbb{R}^{d},\mathbb{R}^{d})$ for all fixed $\nu\in\mathcal{P}(\mathbb{R}^{d})$. Finally, we need to show that
$\nabla D_{\nu}\mathcal{F}^{\eta}(x,\nu)$ $=\int\limits_{\mathbb{R}^{d}}\nabla_{1}b^{\eta}(x,z,\nu)\mathrm{d}\nu(z)=:\nabla b^{1}(x,\nu)+\nabla b^{2}(x,\nu),$
@@
where we used the fact that $\mathrm{k}$ is Lipschitz continuous with Lipschitz constant $\mathtt{M}$ and, for the last term, the dual representation of the $\mathbf{W}_{1}$ distance. In addition, since $\mathrm{k}\in\mathrm{C}^{\infty}(\mathbb{R}^{d}\times\mathbb{R}^{d},[0,+\infty))$ all third derivatives w.r.t. both arguments are bounded, i.e. $\left\|\partial^{3}_{ij\ell}\mathrm{k}(x,y)\right\|\leq\mathtt{M}$, the function
$z\mapsto\lambda\nabla_{2}^{2}\mathrm{k}(z,x_{n})/(\lambda\nu_{n}\left[\mathrm{k}(z,\cdot)\right]+\varphi(z)+\eta)$
@@
for some finite $C$. Since $\lim_{n\to+\infty}(x_{n},\nu_{n})=(x,\nu)$ implies $\left\|x-x_{n}\right\|\to 0$ and $\mathbf{W}_{1}(\nu_{n},\nu)\to 0$, we have that $\nabla b(x,\nu)$ is jointly continuous. Then, the result follows from [32, Theorem 2.11].
Appendix E Particle System and Time Discretization
E.1 Proof of Proposition 5
For $\{x_{1}^{k,N}\}_{k=1}^{N},\{x_{2}^{k,N}\}_{k=1}^{N}\in(\mathbb{R}^{d})^{N}$ and $\ell\in\{1,\dots,N\}$, denote the empirical measures $\pi_{1}^{N}:=(1/N)\sum_{k=1}^{N}\delta_{x_{1}^{k,N}}$ and $\pi_{2}^{N}:=(1/N)\sum_{k=1}^{N}\delta_{x_{2}^{k,N}}$. Using (44) and the fact that for empirical measures
$\mathbf{W}_{1}(\pi_{1}^{N},\pi_{2}^{N})\leq\mathbf{W}_{2}(\pi_{1}^{N},\pi_{2}^{N})\leq N^{-1/2}\left\|x_{1}^{1:N}-x_{2}^{1:N}\right\|\leq\left\|x_{1}^{1:N}-x_{2}^{1:N}\right\|$
where we used the ordering of Wasserstein distances (see, e.g. [57, Remark 6.6]) for the first inequality and the definition of $\mathbf{W}_{2}$ for the second, we have that
@@ E.1 Proof of Proposition 5
The existence and strong uniqueness of a solution to (11) is then a straightforward consequence of the above and [38, Chapter 5, Theorem 2.9 and 2.5].
@@
$\ \ \leq\mathtt{C}_{1}(\left\|x_{1}-x_{2}\right\|+|\pi_{2}\left[\mathrm{k}(x_{2},\cdot)\right]-\pi_{1}\left[\mathrm{k}(x_{1},\cdot)\right]|)+\mathtt{C}_{2}\mathbf{W}_{1}(\pi_{1},\pi_{2})+\alpha\mathtt{L}\left\|x_{1}-x_{2}\right\|$
$\ \ \leq(\mathtt{C}_{1}+\mathtt{C}_{2}+\alpha\mathtt{L})(\left\|x_{1}-x_{2}\right\|+|\pi_{2}\left[\mathrm{k}(x_{2},\cdot)\right]-\pi_{1}\left[\mathrm{k}(x_{1},\cdot)\right]|+\mathbf{W}_{1}(\pi_{1},\pi_{2})|).$
@@
$\qquad\leq\int\limits_{0}^{t}\left\|b(\bm{X}_{s},\pi_{s})-b(\bm{X}_{s}^{1,N},\pi^{N}_{s})\right\|\mathrm{d}s$
@@
$\qquad+(\mathtt{C}_{1}+\mathtt{C}_{2}+\alpha\mathtt{L})\int\limits_{0}^{t}\mathbb{E}\left[\left|\frac{1}{N}\sum_{i=1}^{N}\mathrm{k}(\bm{X}_{s}^{1,N},\bm{X}_{s}^{i,N})-\pi_{s}\left[\mathrm{k}(\bm{X}_{s},\cdot)\right]\right|\right]\mathrm{d}s$
$\qquad+(\mathtt{C}_{1}+\mathtt{C}_{2}+\alpha\mathtt{L})\int\limits_{0}^{t}\mathbb{E}\left[\mathbf{W}_{1}(\pi_{s},\pi_{s}^{N})\right]\mathrm{d}s$
@@
$\qquad+(\mathtt{C}_{1}+\mathtt{C}_{2}+\alpha\mathtt{L})\int\limits_{0}^{t}\mathbb{E}\left[\left|\frac{1}{N}\sum_{i=1}^{N}\mathrm{k}(\bm{X}_{s}^{1,N},\bm{X}_{s}^{i,N})-\pi_{s}\left[\mathrm{k}(\bm{X}_{s},\cdot)\right]\right|\right]\mathrm{d}s,$
where the last inequality follows using the ordering of Wasserstein distances (see, e.g. [57, Remark 6.6]) to obtain $W_{1}(\pi_{s},\pi_{s}^{N})\leq W_{2}(\pi_{s},\pi_{s}^{N})$ and the exchangeability of $\{(\bm{X}_{t}^{k,N})_{t\geq 0}\}_{k=1}^{N}$ to obtain
$\mathbb{E}\left[\mathbf{W}_{2}(\pi_{s},\pi_{s}^{N})\right]$ $\leq\frac{1}{N}\sum_{i=1}^{N}\mathbb{E}\left[\mathbf{W}_{2}(\pi_{s},\delta_{X_{s}^{i,N}})\right]$ (47)
$\leq\frac{1}{N}\sum_{i=1}^{N}\mathbb{E}\left[\left\|\bm{X}_{s}-\bm{X}_{s}^{i,N}\right\|\right]\leq\mathbb{E}\left[\left\|\bm{X}_{s}-\bm{X}_{s}^{1,N}\right\|\right].$
@@
where we used the Lipschitz continuity of $\mathrm{k}$, (47) and the fact that $\{(\bm{X}_{t}^{k,N})_{t\geq 0}\}_{k=1}^{N}$ is exchangeable to obtain the last inequality.
Consider now the last term in the display above and let us denote $\Delta_{i}:=\mathrm{k}(\mathbf{X}_{s}^{1,N},\mathbf{X}_{s}^{i,\star})-\pi_{s}[\mathrm{k}(\mathbf{X}_{s}^{1,N},\cdot)]$, then
$\frac{1}{N}\mathbb{E}\left[\left|\sum_{i=1}^{N}\mathrm{k}(\bm{X}_{s}^{1,N},\bm{X}_{s}^{i,\star})-\pi_{s}\left[\mathrm{k}(\bm{X}_{s},\cdot)\right]\right|\right]$
$\qquad\qquad\leq\frac{1}{N}\mathbb{E}\left[\left|\sum_{i=1}^{N}\Delta_{i}\right|\right]+\mathbb{E}\left[\left|\pi_{s}\left[\mathrm{k}(\bm{X}_{s}^{1,N},\cdot)\right]-\pi_{s}\left[\mathrm{k}(\bm{X}_{s},\cdot)\right]\right|\right]$
$\qquad\qquad\leq\frac{1}{N}\mathbb{E}\left[\left|\sum_{i=1}^{N}\Delta_{i}\right|\right]+\mathtt{M}\mathbb{E}\left[\sup_{u\in[0,s]}\left\|\bm{X}_{u}^{1,N}-\bm{X}_{u}\right\|\right],$
using the Lipschitz continuity of $\mathrm{k}$. For the remaining term, using Jensen’s inequality and the independence of the $\Delta_{i}$s we find
$\frac{1}{N}\mathbb{E}\left[\left|\sum_{i=1}^{N}\Delta_{i}\right|\right]=$ $\frac{1}{N}\mathbb{E}\left[\left\{\left(\sum_{i=1}^{N}\Delta_{i}\right)^{2}\right\}^{1/2}\right]$
$\leq$ $\frac{1}{N}\mathbb{E}\left[\left(\sum_{i=1}^{N}\Delta_{i}\right)^{2}\right]^{1/2}\qquad\textrm{(Jensen)}$
$=$ $\frac{1}{N}\mathbb{E}\left[\left(\sum_{i=1}^{N}\Delta_{i}^{2}+\sum_{i\neq j}\Delta_{i}\Delta_{j}\right)\right]^{1/2}$
$=$ $\frac{1}{N}\mathbb{E}\left[\sum_{i=1}^{N}\Delta_{i}^{2}\right]^{1/2}\qquad\textrm{(Independence)}.$
Plugging the above into (46) we obtain
$\mathbb{E}\left[\sup_{s\in[0,t]}\left\|\bm{X}_{s}-\bm{X}_{s}^{1,N}\right\|\right]$ $\leq 2(1+\mathtt{M})(\mathtt{C}_{1}+\mathtt{C}_{2}+\alpha\mathtt{L})\int\limits_{0}^{t}\mathbb{E}\left[\sup_{u\in[0,s]}\left\|\bm{X}_{u}-\bm{X}_{u}^{1,N}\right\|\right]\mathrm{d}s$
$+\frac{\mathtt{C}_{1}+\mathtt{C}_{2}+\alpha\mathtt{L}}{N}\int_{0}^{t}\mathbb{E}\left[\sum_{i=1}^{N}|\mathrm{k}(\bm{X}_{s}^{1,N},\bm{X}_{s}^{i,\star})-\pi_{s}\left[\mathrm{k}(\bm{X}_{s}^{1,N},\cdot)\right]|^{2}\right]^{1/2}\mathrm{d}s.$
@@
$\mathbb{E}\left[|\mathrm{k}(\bm{X}_{s}^{1,N},\bm{X}_{s}^{i,\star})-\pi_{s}\left[\mathrm{k}(\bm{X}^{1,N}_{s},\cdot)\right]|^{2}\mid\bm{X}_{s}^{1,N}\right]$ $=\textrm{var}\left(\mathrm{k}(\bm{X}_{s}^{1,N},\bm{X}_{s}^{i,\star})\mid\bm{X}_{s}^{1,N}\right)$
@@ E.2 Proof of Proposition 6
with $b^{\eta}$ defined in (10). Recall that under Assumption 1 and 3 we have (44), i.e., there exist $C_{1}\geq 0$ such that for any $x_{1},x_{2}\in\mathbb{R}^{d}$ and $\pi_{1},\pi_{2}\in\mathcal{P}(\mathbb{R}^{d})$
@@
$\leq\lambda\mathtt{M}/\eta+\lambda(\mathtt{M}+1)/\eta=:C_{0}.$
@@
$\qquad\leq 2C_{1}N^{1/2}\left\|x_{1}^{1:N}-x_{2}^{1:N}\right\|,$
where the last line follows from Hölder’s inequality.
@@
$\qquad\leq-\alpha\mathtt{m}\left\|x_{1}^{1:N}-x_{2}^{1:N}\right\|^{2}+\alpha\mathtt{c}N+2C_{0}N^{1/2}\left\|x_{1}^{1:N}-x_{2}^{1:N}\right\|.$
Let $R=\max(8C_{0}N^{1/2}/(\alpha\mathtt{m}),2(\mathtt{c}N/\mathtt{m})^{1/2})$ with $\mathtt{m}$ and $\mathtt{c}$ as in Assumption 3(c). Then
$\alpha\mathtt{c}N\leq\frac{\alpha R^{2}\mathtt{m}}{4}\qquad\qquad 2C_{0}N^{1/2}\leq\frac{\alpha R\mathtt{m}}{4},$
and for any $x^{1:N}_{1},x_{2}^{1:N}\in(\mathbb{R}^{d})^{N}$ with $\left\|x_{1}^{1:N}-x_{2}^{1:N}\right\|\geq R$ we further have
$\langle B_{N}(x^{1:N}_{1})-B_{N}(x^{1:N}_{2}),x_{1}^{1:N}-x_{2}^{1:N}\rangle$ (53)
$\qquad\qquad\leq-\alpha\mathtt{m}\left\|x_{1}^{1:N}-x_{2}^{1:N}\right\|^{2}+\frac{\alpha R^{2}\mathtt{m}}{4}+\frac{\alpha\mathtt{m}R}{4}\left\|x_{1}^{1:N}-x_{2}^{1:N}\right\|$
$\qquad\qquad\leq-(\alpha\mathtt{m}/2)\left\|x_{1}^{1:N}-x_{2}^{1:N}\right\|^{2}.$
We conclude upon combining (51), (52) and [21, Corollary 22] whose assumptions C1 and B3 are satisfied and we verify B2 in (51) which in turns implies B5.
E.3 Proof of Proposition 7
@@
$\bm{M}_{t}^{N}:=\frac{1}{2}\left\|\bm{X}_{t}^{1,N}\right\|^{2}-\int\limits_{0}^{t}\left\{\langle\bm{X}_{u}^{1,N},b(\bm{X}_{u}^{1,N},\pi^{N}_{u})\rangle+(1+\alpha)d\right\}\mathrm{d}u,$
where $\{(\bm{X}_{t}^{k,N})_{t\geq 0}\}_{k=1}^{N}$ is given in (11). Using Itô’s formula we have that, for any $s<t$:
$\frac{1}{2}\left\|\bm{X}_{t}^{1,N}\right\|^{2}$ $=\frac{1}{2}\left\|\bm{X}_{s}^{1,N}\right\|^{2}+\int\limits_{s}^{t}\langle\bm{X}_{u}^{1,N},b(\bm{X}_{u}^{1,N},\pi^{N}_{u})\rangle\mathrm{d}u$
$+(1+\alpha)\int_{s}^{t}\left\|\bm{B}_{u}^{1}\right\|^{2}\mathrm{d}u,$
from which it follows that
$\bm{M}_{t}^{N}$ $=\frac{1}{2}\left\|\bm{X}_{s}^{1,N}\right\|^{2}-\int\limits_{0}^{s}\langle\bm{X}_{u}^{1,N},b(\bm{X}_{u}^{1,N},\pi^{N}_{u})\rangle\mathrm{d}u$
$-(1+\alpha)dt+(1+\alpha)\int_{s}^{t}\left\|\bm{B}_{u}^{1}\right\|^{2}\mathrm{d}u.$
Denote by $(\mathcal{F}_{t}^{N})_{t\geq 0}$ the natural filtration of $\{(\bm{B}_{t}^{k})_{t\geq 0}\}_{k=1}^{N}$. Then,
$\mathbb{E}\left[\bm{M}_{t}^{N}\mid\mathcal{F}_{s}^{N}\right]$ $=\frac{1}{2}\left\|\bm{X}_{s}^{1,N}\right\|^{2}-\int\limits_{0}^{s}\langle\bm{X}_{u}^{1,N},b(\bm{X}_{u}^{1,N},\pi^{N}_{u})\rangle\mathrm{d}u$
$-(1+\alpha)dt+(1+\alpha)d(t-s)$
$=\bm{M}_{s}^{N},$
showing that $\bm{M}_{t}^{N}$ is an $\mathcal{F}^{N}$-local martingale.
@@
$\frac{1}{2}\mathbb{E}\left[\left\|\bm{X}_{t}^{1,N}\right\|^{2}\right]-\frac{1}{2}\mathbb{E}\left[\left\|\bm{X}_{s}^{1,N}\right\|^{2}\right]$
$\qquad=\mathbb{E}\left[\int\limits_{s}^{t}\left\{\langle\bm{X}_{u}^{1,N},b(\bm{X}_{u}^{1,N},\pi^{N}_{u})\rangle+(1+\alpha)d\right\}\mathrm{d}u\right]$
$\qquad=\mathbb{E}\left[\int\limits_{s}^{t}\langle\bm{X}_{u}^{1,N},\int\limits_{\mathbb{R}^{d}}b^{\eta}(\bm{X}_{u}^{1,N},z,\pi^{N}_{u})\mathrm{d}\pi_{u}^{N}(z)\rangle\mathrm{d}u\right]$
$\qquad\phantom{=}-\mathbb{E}\left[\int\limits_{s}^{t}\left\{\langle\bm{X}_{u}^{1,N},\alpha\nabla U(\bm{X}_{u}^{1,N})\rangle-(1+\alpha)d\right\}\mathrm{d}u\right],$
@@
$\frac{1}{2}\mathbb{E}\left[\left\|\bm{X}_{t}^{1,N}\right\|^{2}\right]-\frac{1}{2}\mathbb{E}\left[\left\|\bm{X}_{s}^{1,N}\right\|^{2}\right]$
@@
$\qquad\phantom{=}+\mathbb{E}\left[\int\limits_{s}^{t}\left\{-(1+\alpha)\mathtt{m}\left\|\bm{X}_{u}^{1,N}\right\|^{2}+(1+\alpha)\left\|\nabla U(0)\right\|\left\|\bm{X}_{u}^{1,N}\right\|+(1+\alpha)(\mathtt{c}+d)\right\}\mathrm{d}u\right]$
@@
$\qquad\phantom{=}+\int\limits_{s}^{t}\left\{-(1+\alpha)\mathtt{m}\mathbb{E}\left[\left\|\bm{X}_{u}^{1,N}\right\|^{2}\right]+(1+\alpha)\left\|\nabla U(0)\right\|\mathbb{E}\left[\left\|\bm{X}_{u}^{1,N}\right\|\right]+(1+\alpha)(\mathtt{c}+d)\right\}\mathrm{d}u,$
@@
Let $\mathcal{V}_{t}^{N}=\mathbb{E}\left[\left\|\bm{X}_{t}^{1,N}\right\|^{2}\right]$. Appealing to the fundamental theorem of calculus we get that
$\frac{1}{2}\mathrm{d}\mathcal{V}_{t}^{N}/\mathrm{d}t\leq$ $\mathbb{E}\left[\int\limits_{\mathbb{R}^{d}}\left\|b^{\eta}(\bm{X}_{t}^{1,N},z,\pi^{N}_{t})\right\|\left\|\bm{X}_{t}^{1,N}\right\|\mathrm{d}\pi_{t}^{N}(z)\right]$
$-(1+\alpha)\mathtt{m}\mathbb{E}\left[\left\|\bm{X}_{t}^{1,N}\right\|^{2}\right]+(1+\alpha)\left\|\nabla U(0)\right\|\mathbb{E}\left[\left\|\bm{X}_{t}^{1,N}\right\|\right]$
$+(1+\alpha)(\mathtt{c}+d).$